OpenStudy (anonymous):

MEDAL AND FAN Solve x2 + 8x − 3 = 0 using the completing-the-square method. x = four plus or minus the square root of three x = negative four plus or minus the square root of three x = four plus or minus the square root of nineteen x = negative four plus or minus the square root of nineteen

OpenStudy (anonymous):

x^2+8x-3=0 fits the form ax^2+bx+c where a=1, b=8, and c=-3.

OpenStudy (anonymous):

To complete the form, what do we need to do?

OpenStudy (anonymous):

that's what im having trouble with

OpenStudy (anonymous):

Well, we need to add 16 and subtract 16.

OpenStudy (anonymous):

Want to know where I got 16 from?

OpenStudy (anonymous):

It's really simple.

OpenStudy (anonymous):

Well, 16 is a constant right?

OpenStudy (anonymous):

So we say: Let constant= (b/2)^2

OpenStudy (anonymous):

Then, b=8

OpenStudy (anonymous):

Therefore, K= (b/2)^2= (8/2)^2= 16

OpenStudy (anonymous):

x=-4+/19 the divide thing it goes up then over then the 19 is under it i used mathway.com